OpenStudy (anonymous):

A roast is removed from the freezer at a refrigerator and left on the counter to defrost. The temperature of the roast was -4 degrees centigrade when it was removed from the freezer and t hours later, was inc. at a rate of T'(t) = 7e^.35t. find a formula for the temp of the roast after hours answer is 16-20e^-.35t i dont get where the 16 coming from

OpenStudy (anonymous):

what i have so far \[\int\limits7e^-.35t dt\] \[\int\limits7+ \int\limits e^-.35t\] 7[1/.35e^-.35] my answer -20e^-.35+c

OpenStudy (fibonaccichick666):

ok uhm let's start by correcting your integration\[\int 7e^{-.35t}dt=7\int e^{-.35t}dt\]

OpenStudy (fibonaccichick666):

try again for me please =)

OpenStudy (dan815):

oops i didnt see u had a mission :P

OpenStudy (fibonaccichick666):

haha it's cool

OpenStudy (dan815):

are u there @elskydiablo

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

can you please answer me then, please redo the problem showing your steps

OpenStudy (anonymous):

\[7\int\limits e ^{-.35t } dt\] 7[-1/.35e^-.35] 7/-.35[e^-.35] -20.e^-.35?

OpenStudy (dan815):

ur messing up the integration for e

OpenStudy (anonymous):

how?

OpenStudy (fibonaccichick666):

You are writing it incorrectly but achieving a correct answer

OpenStudy (fibonaccichick666):

can you please write your first step in LaTex

OpenStudy (anonymous):

LaTex?

OpenStudy (fibonaccichick666):

The equation option at the bottom of the dialogue box. It is actually a math script named latex.

OpenStudy (fibonaccichick666):

here copy this to start \[7\int\limits e ^{-.35t } dt\]

OpenStudy (fibonaccichick666):

oops [[7\int\limits e ^{-.35t } dt]]

OpenStudy (fibonaccichick666):

dan change the first and second to last bracket to a \

OpenStudy (dan815):

ok i see

OpenStudy (anonymous):

\[7\int\limits e ^{-.35t} dt\] \[7\int\limits e ^{-.35t}\dt] \[7\int\limits 1/-.35e ^{-.35t}\dt] \[7/-.35 e ^{-.35t}\] is this right so far?

OpenStudy (dan815):

\[7\int\limits e ^{-.35t } dt\] \[=(7/-0.35)e ^{-.35t } + c \] \[=-20e ^{-.35t } + c\]

OpenStudy (dan815):

ya u got it right now, u kept forgetting the t

OpenStudy (dan815):

now apply that inital condition to it

OpenStudy (dan815):

T(0)=-4

OpenStudy (anonymous):

yes i see what i did wrong.

OpenStudy (fibonaccichick666):

yea that's right now you need to find your c, right? and you know that T(0)=-4

OpenStudy (fibonaccichick666):

So what is your next step?

OpenStudy (fibonaccichick666):

in words, you don't have to type it

OpenStudy (anonymous):

t(0)=-20e^-.35(0)+c=-4

OpenStudy (fibonaccichick666):

right next what should you do?

OpenStudy (anonymous):

-20+c=-4 c=16

OpenStudy (fibonaccichick666):

yep! congrats you did it!

OpenStudy (anonymous):

thank you so much

OpenStudy (fibonaccichick666):

np

OpenStudy (dan815):

listen to my elsky! u need to become MORE SERIOUS ABOUT THESE THINGS!

OpenStudy (dan815):

i expect you to be doing boundary value laplace PDE EQUATIONS by the time i see u again

OpenStudy (dan815):

good bye